# DeepCompile: A Compiler-Driven Approach to Optimizing Distributed Deep Learning Training

### v1

@@ 5.2. Efficiency
Figure 7 presents the throughputs observed in our experiments. We used Llama-3-70B with 32 GPUs and Mixtral-8x7B with 16 GPUs. We evaluated DeepSpeed ZeRO-3 both without and with the PyTorch compiler enabled, labeled as ZeRO3 and ZeRO3 (C), respectively. Similarly, we include results for PyTorch FSDP, labeled as FSDP and FSDP (C) depending on whether the PyTorch compiler is enabled. For DeepCompile, we tested three configurations: enabling only proactive prefetching, enabling only selective unsharding, and enabling both in that order. These configurations are labeled as DeepCompile (P), DeepCompile (S), and DeepCompile (P+S), respectively, in the charts. Since the models fit within the available GPU memory at these scales, adaptive offloading was not applied in these experiments.
We conducted this experiment using different batch sizes and sequence lengths. In addition, we evaluated training performance with gradient accumulation.
@@
DeepCompile consistently outperformed all baseline methods, with the most significant gains observed when both proactive prefetching and selective unsharding were enabled (DeepCompile (P+S)). Under the setting with a gradient accumulation step of 1, the smallest improvement over ZeRO-3 was 1.11×, observed in the configuration with sequence length 512 and batch size 1. In this setting, computation is significantly faster relative to communication, leaving little opportunity for our optimizations to hide communication overhead through overlap. As a result, the benefit of prefetching and unsharding becomes limited. For all other conditions, DeepCompile achieved approximately 1.14× to 1.18× speedup over ZeRO-3, and 1.16× to 1.26× over FSDP in the Llama-3-70B experiments. For Mixtral-8x7B, the speedups over ZeRO-3 ranged from 1.07× to 1.35×, while those over FSDP ranged from 1.06× to 1.14×. Mixtral-8x7B is an MoE model, where only a subset of parameters is activated for each input token. Due to its relatively small computation load compared to total parameter size, it suffers more from communication bottlenecks. As a result, similar to Llama-3-70B, the improvements were slightly larger when the batch size and sequence length were increased, enabling better overlap between communication and computation.
@@
Figure 7(a)(iii) and (b)(iii) show throughput results when varying the gradient accumulation steps from 1 to 16. As discussed in Section 4.3, this setting significantly benefits from selective unsharding. In the Llama-3-70B experiment (Fig. 7(a)(iii)), DeepCompile achieves increasingly higher throughput as the accumulation step increases, with up to 1.28× improvement over ZeRO-3 at step 16. In Mixtral-8x7B (Fig. 7(b)(iii)), the improvement is even more pronounced due to the model’s heavier communication load. DeepCompile achieves up to 1.54× higher throughput compared to ZeRO-3 at step 16.
@@ 5.4. Adaptive Offloading
The Llama-3-70B model requires 32 or more GPUs, each with 80GB of memory, to fit entirely into GPU memory under our current settings. However, by offloading optimizer states, we are able to fit the model onto just 16 GPUs. To evaluate the benefits of adaptive offloading of optimizer states, as described in Section 4.4, we compared the throughput achieved using DeepSpeed’s offloading feature with that of our adaptive offloading approach.
Figure 9 shows the iteration times for each method. ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile represent the results of using DeepSpeed ZeRO-3 with and without the PyTorch compiler, respectively. Notably, DeepSpeed not only offloads optimizer states to host memory but also performs parameter updates on the CPU. While this approach saves more memory than DeepCompile’s offloading, it significantly slows down parameter updates. Since this behavior makes it difficult to compare DeepSpeed’s offloading directly with our method—which keeps parameter updates on the GPU—we also included results for DeepCompile (Offload optimizer all + sync), which offloads all optimizer states at the beginning of the forward pass and reloads them synchronously at the end. Our adaptive offloading method is shown as DeepCompile (Offload selective + async), which minimizes the amount of data offloaded and overlaps offloading and reloading asynchronously with computation.
@@ 5.5. Compilation time
As shown in Fig.3, DeepCompile applies a sequence of optimization passes. In our experiments, we first executed the pass that inserts allgather and release operations. After a short warmup period (five training iterations), we applied additional passes including proactive prefetching and selective unsharding.
@@
The variation in compile time across optimization configurations for the same model is less than 10%, indicating that the additional cost of enabling proactive prefetching or selective unsharding is relatively minor compared to the base analysis and scheduling pass. On the other hand, compile times vary more significantly across models, as seen in the longer durations for Mixtral-8x7B compared to Llama-3-70B. This difference stems from model-specific factors such as graph size and layer composition.
@@ 5.6. Correctness
To verify the correctness of the optimization passes in DeepCompile, we compared the resulting loss values with those of DeepSpeed ZeRO-3. We initialized Llama-3 70B model with random weights but the same random seed in both settings. We used the AG News corpus (Zhang et al., 2016) as training examples, with the sequence length and micro batch size set to 512 and 1, respectively. The learning rate was set to 1.5e-5. Figure 10 shows the training losses for both settings. Although some operators are non-deterministic and introduce subtle differences, the loss curves were closely aligned.
@@ 6. Related Work
6.1. Distributed Training Framework
As discussed in Section 2, various distributed training strategies have been proposed to scale deep learning models beyond the capacity of a single GPU. For instance, tensor parallelism, as introduced in Megatron-LM (Shoeybi et al., 2019), splits large matrix operations across devices, while pipeline parallelism, as implemented in GPipe (Huang et al., 2019), assigns consecutive layers to different GPUs. While effective, both approaches require the user to manually modify the model to specify how tensors or layers are partitioned. This makes it challenging to adapt these strategies to diverse or rapidly evolving model architectures.
In contrast, our work focuses on the fully sharded approach, as implemented in ZeRO (Rajbhandari et al., 2020) and FSDP (Team, 2023), which requires no model code changes and is applicable to a wide range of architectures. DeepCompile builds on this fully sharded approach and introduces compiler-level optimizations to improve its efficiency.
Note that DeepCompile is not inherently limited to the fully sharded approach. Its design is compatible with other parallelization strategies. For example, RaNNC (Tanaka et al., 2021) automatically partitions the computation graph to enable pipeline parallelism, and Alpa (Mo et al., 2022) selects distributed execution rules based on the partitioned state of tensors. Similarly, DeepCompile could support automated parallelization by replacing computation operators in the graph and integrating this with other optimization passes within a unified framework.
6.2. Optimizing Parallelization Strategies
To improve training efficiency for large-scale models, various strategies such as data, tensor, and pipeline parallelism have been studied. Recent work has focused on automatically discovering optimal combinations of these strategies, including Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), Unity (Unger et al., 2022), and Galvatron (Miao et al., 2022). Dynamic switching techniques have also been proposed, such as HotSpa (Ge et al., 2024), which aims to adaptively reconfigure parallelization strategies during training.
However, these approaches generally assume that the model is already implemented in a way that supports the desired forms of parallelism. In practice, tensor and pipeline parallelism often require model-specific partitioning logic and integration with low-level APIs. Supporting multiple strategies and enabling dynamic switching can involve considerable engineering effort, particularly for complex or evolving model architectures. As a result, despite their promise, such automated parallelization systems have seen limited adoption in current practice.
6.3. Computation-Communication Overlapping
As discussed earlier, widely used frameworks such as ZeRO-3 and FSDP incorporate prefetching techniques to overlap communication with computation and reduce latency from all-gather operations. Prefetching has become a standard optimization for improving efficiency in fully sharded training.
More recently, techniques for overlapping communication at a finer granularity, such as intra-layer parallelism, have also been proposed (Wang et al., 2022, 2024). These approaches aim to eliminate or hide communication costs by restructuring computation within layers. In contrast, our work focuses on profiling-guided, graph-level optimization of communication overlap.
Like DeepCompile, SimpleFSDP (Zhang et al., 2024) targets compiler-level optimization of the fully sharded approach. It performs memory-aware fusion of all-gather operations to reduce communication overhead, but does not take into account dynamic changes in memory usage when deciding how early each all-gather can be issued without exceeding memory constraints. Moreover, SimpleFSDP focuses solely on scheduling communication for FSDP. In contrast, DeepCompile is designed to support a broader range of optimizations and enables techniques such as selective unsharding and adaptive offloading.
@@ 7. Conclusion
We introduced DeepCompile, a compiler-based framework for optimizing distributed training of large-scale deep learning models. DeepCompile starts from a standard single-GPU computation graph and incrementally inserts distributed operators such as all-gather and release, enabling graph-level analysis and holistic optimization of distributed training. Unlike existing frameworks like ZeRO-3 and FSDP, which rely on heuristics and fixed module boundaries, DeepCompile leverages runtime profiling to flexibly reorder and coordinate operations based on actual memory usage and data dependencies.
Built on this foundation, DeepCompile implements a fully sharded training approach along with three key optimizations: proactive prefetching to improve communication–computation overlap, selective unsharding to reduce redundant communication, and adaptive offloading to dynamically manage memory under tight resource constraints. These optimization passes are applied in sequence, informed by profiling feedback, allowing DeepCompile to coordinate them effectively in a unified compilation framework.
Our experiments demonstrate that DeepCompile consistently outperforms ZeRO-3 and FSDP baselines across a range of configurations. It achieves up to 1.28× and 1.54× higher throughput for Llama-3 70B and Mixtral 8×7B models, respectively. Moreover, in scenarios with limited GPU memory, DeepCompile’s adaptive offloading achieves up to a 7.01× throughput improvement by automatically overlapping communication and reducing memory pressure.
DeepCompile represents a step forward in compiler-driven distributed training. Its graph transformation capabilities and runtime-aware optimization passes offer a scalable and automated path toward efficient parallelization. In future work, we aim to extend DeepCompile to support broader classes of optimizations, such as fully automated parallelization planning and enhanced memory scheduling.

### v2

@@ 5.2. Efficiency
Figure 7 presents the throughputs observed in our experiments. We used Llama-3-70B and Mixtral-8x7B with 32 GPUs. All throughput numbers reflect end-to-end time per training iteration, including the forward and backward passes as well as parameter updates. We evaluated ZeRO-3 both without and with the PyTorch compiler enabled, labeled as ZeRO3 and ZeRO3 (C), respectively. Similarly, we include results for FSDP, labeled as FSDP and FSDP (C) depending on whether the PyTorch compiler is enabled. For DeepCompile, we tested three configurations: enabling only proactive prefetching, enabling only selective unsharding, and enabling both in that order. These configurations are labeled as DeepCompile (P), DeepCompile (S), and DeepCompile (P+S), respectively, in the charts. Since the models fit within the available GPU memory at these scales, adaptive offloading was not applied in these experiments.
We conducted experiments with varying batch sizes, sequence lengths, and gradient accumulation steps. The per-GPU batch size (number of sequences) was set to 1, 2, or 4, and the sequence length to 512, 1024, or 2048. We also varied the gradient accumulation steps from 1 to 16. The effective batch size, measured in tokens, is given by: $\text{\#Sequences per GPU}\times\text{Sequence length}\times\text{Accumulation steps}\times\text{\#GPUs}$. Under these settings, the total number of tokens per step ranged from 8K to 4M, covering a broad range representative of common configurations used for training models at this scale.
For the system parameters defined in Table 1, we set $M$ to 90% of the available GPU memory to maintain a safety margin. $M_{\text{prefetch}}$ was set to 2 GB. Increasing this value allows DeepCompile to schedule more all-gather communications earlier, but we did not observe any clear performance benefit from setting it higher. The parameter $\alpha$ controls the aggressiveness of all-gather fusion. While larger values encourage more aggressive fusion, we found that increasing $\alpha$ beyond 1.5 had little effect on performance. Based on experiments with values ranging from 1.0 to 2.0, we fixed $\alpha$ to 1.5. We used these values throughout all of the experiments.
@@
DeepCompile consistently outperformed all baselines, with the largest gains observed when both proactive prefetching and selective unsharding were enabled (DeepCompile (P+S)). With gradient accumulation set to 1, the smallest improvement over ZeRO-3 for Llama-3-70B (Fig 7(a),(i)(ii)) was 1.11×, observed at sequence length 512 and batch size 1. In this configuration, computation dominates communication, leaving limited room for our optimizations to hide communication latency via overlap. Consequently, the benefits of prefetching and unsharding are reduced. Under all other conditions, DeepCompile achieved a 1.14×–1.18× speedup over ZeRO-3 and 1.16×–1.26× over FSDP.
For Mixtral-8x7B, the speedups over ZeRO-3 ranged from 1.07× to 1.35×, while those over FSDP ranged from 1.06× to 1.14× (Figure 7(b),(i)(ii)). Mixtral-8x7B is an MoE model, where only a subset of parameters is activated for each input token. Due to its relatively small computation load compared to total parameter size, it suffers more from communication bottlenecks. As a result, similar to Llama-3-70B, the improvements were slightly larger when the batch size and sequence length were increased, enabling better overlap between communication and computation.
@@
Figure 7(a)(iii) and (b)(iii) show throughput results when varying the gradient accumulation steps from 1 to 16. All numbers reflect the end-to-end time per training iteration, including all gradient accumulation steps. As discussed in Section 4.3, this setting significantly benefits from selective unsharding. In the Llama-3-70B experiment (Fig. 7(a)(iii)), DeepCompile achieves increasingly higher throughput as the accumulation step increases, with up to 1.28× improvement over ZeRO-3 at step 16. In Mixtral-8x7B (Fig. 7(b)(iii)), the improvement is even more pronounced due to the model’s heavier communication load. DeepCompile achieves up to 1.54× higher throughput compared to ZeRO-3 at step 16.
@@ 5.4. Adaptive Offloading
The Llama-3-70B model requires at least 32 GPUs (each with 80GB memory) to fit entirely into GPU memory under our settings. However, by offloading optimizer states, we are able to run the model on just 16 GPUs. To evaluate the benefits of adaptive optimizer offloading (Section 4.4), we compared the throughput achieved by DeepSpeed’s offloading mechanism combined with ZeRO-3 against our adaptive offloading approach.
Figure 9 presents the end-to-end iteration time for each method. ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile represent the results of using ZeRO-3 with and without the PyTorch compiler, respectively. Notably, ZeRO-3 not only offloads optimizer states to host memory but also performs parameter updates on the CPU. While this approach saves more memory than DeepCompile ’s offloading, it significantly slows down parameter updates. Since this behavior makes it difficult to compare ZeRO-3’s offloading directly with our method, which keeps parameter updates on the GPU, we also included results for DeepCompile (Offload optimizer all + sync), which offloads all optimizer states at the beginning of the forward pass and reloads them synchronously at the end. Our adaptive offloading method is shown as DeepCompile (Offload selective + async), which minimizes the amount of data offloaded and overlaps offloading and reloading asynchronously with computation.
@@ 5.5. Compilation time
As shown in Fig.3, DeepCompile applies a sequence of optimization passes. In our experiments, we first executed the pass that inserts allgather and release operations. After a short warmup period (five training iterations in out experiments), we applied additional passes including proactive prefetching and selective unsharding.
@@
The variation in compile time across optimization configurations for the same model is less than 10%, indicating that the additional cost of enabling proactive prefetching or selective unsharding is relatively minor compared to the base analysis and the pass to enable the fully-shareded approach. On the other hand, compile times vary more significantly across models, as seen in the longer durations for Mixtral-8x7B compared to Llama-3-70B. This difference stems from model-specific factors such as graph size and layer composition.
@@ 5.6. Correctness
To verify the correctness of the optimization passes in DeepCompile, we compared the resulting loss values with those of ZeRO-3. We initialized Llama-3 70B model with random weights but the same random seed in both settings. We used the AG News corpus (Zhang et al., 2016) as training examples, with the sequence length and micro batch size set to 512 and 1, respectively. The learning rate was set to 1.5e-5. Figure 10 shows the training losses for both settings. Although some operators are non-deterministic and introduce subtle differences, the loss curves were closely aligned.
@@
Table 3. Comparison of DeepCompile with existing distributed training systems
Feature ZeRO-3 / FSDP (Rajbhandari et al., 2020; Team, 2023) SimpleFSDP (Zhang et al., 2024) Alpa / Unity (Mo et al., 2022; Unger et al., 2022) DeepCompile (ours)
No code modification
Compiler-level optimization
Profiling guided optimization
Communication overlap
Composable optimizations
Memory-aware scheduling
@@ 6. Related Work
Distributed training has become a central topic in deep learning system design, leading to a range of frameworks that support model parallelism, memory optimization, and scaling strategies. We summarize key systems and contrast their capabilities with our approach in Table 3.
Framework-level sharding Fully sharded training frameworks such as ZeRO-3 (Rajbhandari et al., 2020) and PyTorch FSDP (Team, 2023) partition model parameters and optimizer states across devices and apply runtime optimizations like prefetching and unsharding. These optimizations are typically implemented using hooks inserted at the framework level and rely on heuristics that do not leverage global analysis or profiling feedback. As a result, they offer limited control over the scheduling of communication and memory operations, and cannot coordinate multiple interacting optimizations in a globally informed manner.
Compiler-based prefetching SimpleFSDP (Zhang et al., 2024) uses a compiler to improve communication scheduling for fully sharded training. It analyzes operator-level memory usage to decide whether to fuse multiple all-gather operations. However, it focuses solely on prefetching and does not support coordination across multiple optimizations or feedback loops between passes. In addition, while it considers memory usage for fusion decisions, it does not take memory dynamics into account when determining the placement of all-gather operations. DeepCompile generalizes this idea by introducing a pass-based architecture in which each optimization pass can access updated profiling results and respond to memory dynamics.
Automatic parallelization planning FlexFlow (Jia et al., 2019) proposed selecting optimal combinations of data, tensor, and pipeline parallelism based on cost models. Building on this idea, recent systems such as Alpa (Mo et al., 2022), nnScaler (Lin et al., 2024), and Unity (Unger et al., 2022) integrate parallelism planning with compiler-based optimization techniques. These systems focus on cost-model-driven planning and static graph analysis to combine parallelism strategies and GPU-local optimizations. However, they do not account for runtime memory usage trends or support techniques such as communication–computation overlap and memory-aware offloading.
DeepCompile complements these approaches by emphasizing profiling-guided, memory-centric optimization using modular passes that adapt to runtime behavior. Although our current focus is on the fully sharded approach, DeepCompile can also express other parallelism strategies. For example, tensor parallelism can be realized by identifying Transformer blocks and injecting all-reduce operations; pipeline parallelism can be enabled by partitioning the graph into stages with balanced compute and memory load; and sequence parallelism (Jacobs et al., 2023; Liu et al., 2023) can be supported through all-to-all or peer-to-peer communication.
A key distinction is that DeepCompile’s modular pass framework enables coordinated composition of such parallelization strategies with memory-aware techniques like offloading. This interaction has received little attention in prior work and represents a promising direction for future research.
@@ 7. Conclusion
We presented DeepCompile, a compiler-based optimization framework for distributed deep training. By transforming user-defined models into computation graphs and applying a series of profiling-guided optimization passes, DeepCompile enables fine-grained memory-aware scheduling and communication planning across passes. This approach contrasts with existing systems like ZeRO-3 and FSDP, which lack flexibility in communication timing and memory management.
Built on a fully sharded foundation, DeepCompile incorporates three key optimization passes: proactive prefetching to improve communication–computation overlap, selective unsharding to reduce communication, and adaptive offloading to manage memory when resources are limited. Each pass utilizes updated profiling feedback, allowing the system to compose optimizations based on their combined effects.
Our evaluation demonstrates that DeepCompile consistently outperforms existing baselines. It achieves up to 1.28× and 1.54× throughput improvements on Llama 3 70B and Mixtral 8×7B models, respectively. In memory-constrained settings, adaptive offloading enables up to 7.01× throughput gains by overlapping data transfers with computation.
DeepCompile represents a step forward in compiler-driven distributed training. Its graph transformation capabilities and runtime-aware optimization passes provide a scalable and automated path to efficient parallelization. As discussed in Section 6, future work will extend DeepCompile to support a broader class of optimizations, including communication and memory scheduling.